OpenStudy (anonymous):

After a discount of 10%, the price of a pair of jeans, including tax, was $69.62. The following equation can be used to find x, the price of the pair of jeans without the discount. x - 0.1x + 0.07(x - 0.1x) = 69.62 What is the price of the pair of jeans without the discount?

OpenStudy (anonymous):

Just solve it algebraically. x - 0.1x + 0.07(x - 0.1x) = 69.62 x - 0.1x + 0.07x - 0.007x = 69.62 0.9x + 0.063x = 69.62 0.963x = 69.62 x ≈ 72.29